[figure: 9.7 cm]

DIAGONALIZATION OF A BOSONIC QUADRATIC FORM USING CCM: APPLICATION ON A SYSTEM WITH TWO INTERPENETRATING SQUARE LATTICE ANTIFERROMAGNETS

Sven E. Krüger, Dirk Schmalfuß, Johannes Richter Institut für Theoretische Physik Universität Magdeburg PF 4120, D-39016 Magdeburg, Germany

While the diagonalization of a quadratic bosonic form can always be done using a Bogoliubov transformation, the practical implementation for systems with a large number of different bosons is a tedious analytical task. Here we use the coupled cluster method (CCM) to exactly diagonalise such complicated quadratic forms. This yields to a straightforward algorithm which can easily be implemented using computer algebra even for a large number of different bosons. We apply this method on a Heisenberg system with two interpenetrating square lattice antiferromagnets, which is a model for the quasi 2D antiferromagnet Ba_{2}Cu_{3}O_{4}Cl_{2}. Using a four-magnon spin wave approximation we get a complicated Hamiltonian with four different bosons, which is treated with CCM. Results are presented for magnetic ground state correlations.

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : poster

Corresponding author : Sven E. Krüger

Address for correspondence : Institut für Theoretische Physik Universität Magdeburg PF 4120, D-39016 Magdeburg, Germany

Email address : sven.krueger@physik.uni-magdeburg.de